# footnotetext: ${}^{\text{\Letter}}$ Corresponding Authors

### v1

@@ IV.1 Joule-Thomson coefficient
To study cooling/heating phase of the thermodynamic system, the Joule-Thomson coefficient $\mu$ is introduced, and is defined by the change of the temperature with respect to pressure under a fixed enthalpy given by Okcu:2016tgt ; Feng:2020swq
@@
Use the Eqs.(34), (39) and (18), we obtain the enthalpy of $n$-dimensional FRW universe is given by
@@
see the related discussion about negativity/positivity condition of enthalpy in Appendix B. From (43), one writes the temperature as a function of ${\cal H}$ and $R_{A}$ as follows
@@
and then substituting the Eq.(44) into (39), the pressure $P$ can be rewritten as a function of ${\cal H}$ and $R_{A}$,
@@
$=$ $\frac{4(n-2)\pi^{n/2}R_{A}^{n+3}(R_{A}^{2}+2\widetilde{\alpha})^{2}+16(n-1)\Gamma[(n-1)/2]\pi^{3/2}{\cal H}R_{A}^{8}[(n-2)(R_{A}^{2}+2(n+2)\widetilde{\alpha})]}{(n-2)(n-1)(R_{A}^{2}+2\widetilde{\alpha})^{2}[4(n-1)\Gamma[(n-1)/2]\pi^{3/2}{\cal H}R_{A}^{5}+(n-2)\pi^{n/2}R_{A}^{n}(R_{A}^{2}+2\widetilde{\alpha})]}$
@@ IV.2 Inversion temperature and inversion pressure
In the absence of the Gauss-Bonnet term with $\alpha=0$ (or $\widetilde{\alpha}=0$), we can derive the Joule-Thomson coefficient $\mu$ for an $n$-dimensional FRW universe in Einstein gravity from equations (46) and (47). In this scenario, a positive value of enthalpy (${\cal H}>0$) leads to results indicating that $\mu>0$. This implies that the FRW universe with arbitrary dimensions is always in a cooling stage, indicating the absence of an inversion temperature and inversion pressure. By setting $\mu=0$, we obtain the enthalpy for the existence of an inversion point given by
@@
of the $n$-dimensional FRW universe in Einstein gravity for the negative enthalpy, $i.e.$ ${\cal H}<0$. The inversion point corresponds to a constraint on perfect fluid, which derivations and physical interpretations are written in the Appendix C. From Eqs.(49) and (50), we also obtain the simple relation between the inversion temperature and the inversion pressure
@@
We observe from Fig.1 that, the inversion temperature rises as the inversion pressure increases. However, this rise becomes less significant as the spacetime dimension $n$ increases, indicating a decreasing rate of change in the inversion temperature with respect to the inversion pressure.
@@
${\cal H}=-\frac{(n-2)\pi^{(n-3)/2}R_{A}^{n-5}(R_{A}^{2}+2\widetilde{\alpha})^{2}}{4\Gamma[(n-1)/2](n-1)[(n-2)(R_{A}^{2}+2(n+2)\widetilde{\alpha})]}\,.$ (52)
The enthalpy ${\cal H}<0$ for $\widetilde{\alpha}>0$, and relate this with the result given in (66) one can confirm that there is an inversion point. For $\widetilde{\alpha}<0$, relating Eq.(52) with (67) and (70) we see that: when $R_{A}>\sqrt{-2(n+2)\widetilde{\alpha}}$ the enthalpy ${\cal H}<0$; when $\sqrt{-2\widetilde{\alpha}}<R_{A}<\sqrt{-2(n+2)\widetilde{\alpha}}$ the enthalpy ${\cal H}>0$. Setting $\mu=0$, we also obtain the inversion temperature from Eq. (47) as
$T_{i}=\frac{(n-3)(R_{A}^{2}+2\widetilde{\alpha})-4\widetilde{\alpha}}{2(n-2)\pi R_{A}[R_{A}^{2}+2(n+2)\widetilde{\alpha}]}\,.$ (53)
@@
$P_{i}=\frac{(n-2)\{n(n-3)R_{A}^{4}+\widetilde{\alpha}[n(3n-13)+2]R_{A}^{2}+2n(n-5)\widetilde{\alpha}^{2}\}}{16(n-2)\pi R_{A}^{4}[R_{A}^{2}+2(n+2)\widetilde{\alpha}]}\,.$ (54)
@@
We can observe from Fig.2 that the inversion temperature of the FRW universe exhibits distinct behaviors depending on whether $\widetilde{\alpha}<0$ or $\widetilde{\alpha}>0$ as the inversion pressure varies. Specifically, as shown in Fig.2(a) and Fig.2(b), the inversion temperature increases as the inversion pressure rises for $\widetilde{\alpha}<0$. This increase becomes less significant as the spacetime dimension increases, indicating a decreasing rate of change. In addition, the inversion temperature also increases as the value of the coupling parameter rises. On the other hand, as shown in Fig.2(c) and Fig.2(d), the behavior of the inversion temperature is influenced by the spacetime dimension for $\widetilde{\alpha}>0$. In this case, the inversion temperature decreases as the coupling parameter increases, contrasting with the behavior observed for $\widetilde{\alpha}<0$, where the coupling parameter exerted an opposing influence on the inversion temperature.
@@ IV.3 Isenthalpic curves
The Joule-Thomson expansion is an isenthalpic process. Therefore, it is interesting to study the isenthalpic curves of the FRW universe. In the followings, we plot the inversion and isenthalpic curves of the FRW universe in $T$-$P$ plane with fixed ${\cal H}$ and $\widetilde{\alpha}$, and determine the cooling-heating region.
Refer to caption (a)Refer to caption(b)
Refer to caption (c)Refer to caption(d)
Figure 3: Einstein gravity: The inversion and isenthalpic curves of the FRW universe with dimensions $n=4,5,6,7$. The black line represents the inversion curve, and the rainbow curves intersect with it at points where they exhibit maxima. The values of ${\cal H}$ are marked on $T$-$P$ plane.
@@
Refer to caption (a)Refer to caption(b)
Refer to caption (c)Refer to caption(d)
Refer to caption (e)Refer to caption(f)
Figure 4: Einstein-Gauss-Bonnet gravity: The inversion and isenthalpic curves of the FRW universe with dimensions $n=5,6,7$. The black line represents the inversion curve, and the rainbow curves intersect with it at points where they exhibit maxima. The values of ${\cal H}$ and $\widetilde{\alpha}$ are marked on $T$-$P$ plane.
It can be seen from Fig.4 that under Einstein-Gauss-Bonnet gravity, the inversion point of the FRW universe increases with the enthalpy and spacetime dimension. Through careful analysis, we also find that when the value of the gravitational coupling parameter approaches 00, the behaviour of isenthalpic curves of the FRW universe is similar to that of Einstein’s gravity. When the values of the coupling parameter deviate from 00, the behavior of the isenthalpic curves of the FRW universe is obviously different from the Einstein gravity case. It is evident that the maximum value of the isenthalpic curve of the of the FRW universe increases with both enthalpy and spcatime dimension.
@@ V Conclusions and Discussions
In this paper, we have studied the thermodynamics especially the equation of state and Joule-Thomson expansion of the $n$-dimensional FRW universe with a perfect fluid. From the negative surface gravity of the FRW universe on apparent horizon, $i.e.$ $\kappa<0$, we imposed a constraint on a perfect fluid in the framework of Einstein gravity and Einstein-Gauss-Bonnet gravity respectively, as given in Eqs.(60) and (36). We have derived the equations of state for the $n$-dimensional FRW universe, where the thermodynamic pressure $P$ is defined by the work density $W$, which is a natural definition directly read out from the first law of thermodynamics. We find that the $n$-dimensional FRW universe under Einstein’s gravity with perfect fluid do not have $P$-$V$ phase transitions and critical behavior. Moreover, we find that the specific capacity of the $n$-dimensional FRW universe is always positive. This result is remarkable, since it implies that the FRW universe is stable. Under Einstein-Gauss-Bonnet gravity, the appearance of $P$-$V$ phase transitions as well as (un)stability of the FRW universe with a perfect fluid are depended on the spacetime dimension $n$ and gravitational coupling parameter $\widetilde{\alpha}$.
Furthermore, we present the Joule-Thomson expansion as an application of thermodynamic equations of state to explain the thermodynamic properties of the $n$-dimensional FRW universe. We derived the Joule-Thomson coefficient, and find that the equation of state from Einstein gravity shows that there is an inversion point for the FRW universe with arbitrary dimensions if its enthalpy ${\cal H}$ is negative. If the enthalpy ${\cal H}$ is positive (${\cal H}>0$), the FRW universe with arbitrary dimensions is always in the cooling stage, $i.e.$ there is no inversion temperature and inversion pressure. As an extension, we also calculate the inversion temperature and inversion pressure for FRW universe in Einstein-Gauss-Bonnet gravity with arbitrary dimensions, and illustrate the characteristics of inversion curves and isenthalpic curves in the $T$-$P$ plane. We find that the appearance of an inversion point depends on the spacetime dimension $n$ and gravitational coupling parameter $\widetilde{\alpha}$. Moreover, we also find that the characteristics of inversion temperature exhibits distinct behaviors depending on whether $\widetilde{\alpha}<0$ or $\widetilde{\alpha}>0$ as the inversion pressure varies.
It is worth mentioning that, in the present work we introduce the $T$-$P$ isenthalpic expansion process in a high-dimensional FRW universe within the context of modified gravity theory, pioneering the exploration of thermodynamic properties in arbitrary dimensions for the first time. This results may provides theoretical platform for future observational evidences from astronomy, and it may deepen our understanding the evolution of our universe in the view of thermodynamics. It would also be interesting to explore the same study reported here to higher order modified gravity theories such as Lovelock gravity, they may reveal intriguing phenomena. We will carry these investigations in the future. Moreover, as discussed in Appendix C, we examined the constraints on the perfect fluid of the FRW universe derived from the cooling-heating transition point, which is consistent with the constraint from negative surface gravity, as discussed in (60). These results are related to many other interesting problems, and lead to more physical insights into the understanding of the evolution of the FRW universe from the viewpoint of thermodynamics.
@@ VI Acknowledgment
We thank Profs.Ya-Peng Hu and Yu-Sen An for useful discussions and valuable suggestions that helped us to improve the quality of the paper significantly. This work is supported by the Kashi University high-level talent research start-up fund project under Grant No. 022024002, and the Natural Science Foundation of Sichuan Province under Grant No. 24NSFSC1618.
@@ Appendix A Negativity Condition of Surface Gravity for the Arbitrary Dimensional FRW Universe
In this appendix, we give the negative condition for the surface gravity (6) in Einstein gravity and Einstein-Gauss-Bonnet gravity, respectively.
Since
$\frac{\dot{R}_{A}}{2HR_{A}}<1\,,$ (55)
we can substitute Eqs. (4) and (5) into (55) to see that $\kappa|_{R=R_{A}}<0$ is equivalent to
$-\frac{R_{A}^{2}}{2}\Big{(}\dot{H}-\frac{k}{a^{2}}\Big{)}<1\,,$ (56)
which, in turn, using Eqs. (29) and (30), leads to
$\frac{4\pi(\rho+p)}{n-2}\frac{R_{A}^{2}}{1+{2\widetilde{\alpha}}/{R_{A}^{2}}}<1\,.$ (57)
Thus, we have
$\rho+p<\frac{n-2}{4\pi R_{A}^{2}}\Big{(}1+\frac{2\widetilde{\alpha}}{R_{A}^{2}}\Big{)}\,.$ (58)
Rearranging (58) with the second Friedmann’s equation (30) and Eq.(4), we obtain the negative condition of surface gravity in Einstein-Gauss-Bonnet gravity
$1+\frac{p}{\rho}<\frac{4}{n-1}\Big{(}1+\frac{\widetilde{\alpha}}{R_{A}^{2}+\widetilde{\alpha}}\Big{)}\,,$ (59)
as in Eq. (36). In the limit $\alpha\rightarrow 0$, the condition equation (59) becomes
$p+\frac{n-5}{n-1}\rho<0\,,$ (60)
which is the result of negative surface gravity condition for the $n$-dimensional FRW universe in Einstein gravity.

### v2

@@ IV.1 Joule-Thomson coefficient
To study cooling/heating properties of the thermodynamic system, the Joule-Thomson coefficient $\mu$ is introduced, and is defined by the change of the temperature with respect to pressure under a fixed enthalpy given by Okcu:2016tgt ; Feng:2020swq
@@
Use the Eqs.(34), (39) and (18), we obtain the enthalpy of $n$-dimensional FRW universe is given by [^7] It can be directly deduced that the condition ${\cal H}<0$ is equivalent to $\rho+p>0$. This condition is consistent with the Null Energy Condition (NEC)- a fundamental physical constraint in cosmology, which is generally satisfied by most matter fields.
@@
[…]
@@
[…]
@@
$=$ $\frac{4(n-2)\pi^{n/2}R_{A}^{n+3}(R_{A}^{2}+2\widetilde{\alpha})^{2}+16(n-1)\Gamma[(n-1)/2]\pi^{3/2}{\cal H}R_{A}^{8}[(n-2)R_{A}^{2}+2(n-4)\widetilde{\alpha}]}{(n-2)(n-1)(R_{A}^{2}+2\widetilde{\alpha})^{2}[4(n-1)\Gamma[(n-1)/2]\pi^{3/2}{\cal H}R_{A}^{5}+(n-2)\pi^{n/2}R_{A}^{n}(R_{A}^{2}+2\widetilde{\alpha})]}$
@@ IV.2 Inversion temperature and inversion pressure
In the absence of the Gauss-Bonnet term with $\alpha=0$ (or $\widetilde{\alpha}=0$), we can derive the Joule-Thomson coefficient $\mu$ for an $n$-dimensional FRW universe in Einstein gravity from equations (46) and (47). In this scenario, a positive enthalpy (${\cal H}>0$) leads to $\mu>0$, indicating that the FRW universe in arbitrary dimensions is always cooling, without inversion temperature or pressure. By setting $\mu=0$, we obtain the enthalpy for the existence of an inversion point given by
@@
of the $n$-dimensional FRW universe in Einstein gravity for the negative enthalpy, $i.e.$ ${\cal H}<0$. From Eqs.(49) and (50), we also obtain the simple relation between the inversion temperature and the inversion pressure
@@
We observe from Fig.1 that, the inversion temperature rises as the inversion pressure increases. However, this rise becomes less slowly as the spacetime dimension $n$ increases, indicating a decreasing rate of change in the inversion temperature with respect to the inversion pressure.
@@
${\cal H}=-\frac{(n-2)\pi^{(n-3)/2}R_{A}^{n-5}(R_{A}^{2}+2\widetilde{\alpha})^{2}}{4\Gamma[(n-1)/2](n-1)[(n-2)R_{A}^{2}+2(n-4)\widetilde{\alpha}]}\,$. (52)
[…]
Setting $\mu=0$, we also obtain the inversion temperature from Eq.(47) as
$T_{i}=\frac{(n-3)(R_{A}^{2}+2\widetilde{\alpha})-4\widetilde{\alpha}}{2\pi R_{A}[(n-2)R_{A}^{2}+2(n-4)\widetilde{\alpha}]}\,$. (53)
@@
$P_{i}=\frac{(n-2)\{n(n-3)R_{A}^{4}+\widetilde{\alpha}[n(3n-13)+2]R_{A}^{2}+2n(n-5)\widetilde{\alpha}^{2}\}}{16\pi R_{A}^{4}[(n-2)R_{A}^{2}+2(n-4)\widetilde{\alpha}]}\,$. (54)
@@
We can observe from Fig.2 that the inversion temperature of the FRW universe exhibits distinct behaviors depending on whether $\widetilde{\alpha}<0$ or $\widetilde{\alpha}>0$ as the inversion pressure varies. Specifically, as shown in Fig.2(a) and Fig.2(b), the inversion temperature increases as the inversion pressure rises for $\widetilde{\alpha}<0$. This increase becomes less slowly as the spacetime dimension increases, indicating a decreasing rate of change. In addition, the inversion temperature also increases as the value of the coupling parameter rises. On the other hand, as shown in Fig.2(c) and Fig.2(d), the behavior of the inversion temperature is influenced by the spacetime dimension for $\widetilde{\alpha}>0$. In this case, the inversion temperature decreases as the coupling parameter increases, contrasting with the behavior observed for $\widetilde{\alpha}<0$, where the coupling parameter exerted an opposing influence on the inversion temperature.
@@ IV.3 Isenthalpic curves
[…]
(a)Refer to caption(b)
(c)Refer to caption(d)
[…]
@@
(a)Refer to caption(b)
(c)Refer to caption(d)
(e)Refer to caption(f)
[…]
It can be seen from Fig.4 that under Einstein-Gauss-Bonnet gravity, the inversion point of the FRW universe increases with the enthalpy and spacetime dimension. Through careful analysis, we also find that when the value of the gravitational coupling parameter approaches 0, the behaviour of isenthalpic curves of the FRW universe is similar to that of Einstein’s gravity. When the values of the coupling parameter deviate from 0, the behavior of the isenthalpic curves of the FRW universe is obviously different from the Einstein gravity case. It is evident that the maximum value of the isenthalpic curve of the FRW universe increases with both enthalpy and spcatime dimension.
IV.4 Constraints on the Perfect Fluid in the FRW Universe from the Inversion Point ($\mu=0$)
In what follows, we derive constraints on the perfect fluid of the FRW universe from the inversion point ($\mu=0$).
In the case where $\mathscr{H}<0$, the existence of an inversion point–corresponding to a peak in the $T$-$P$ diagram during Joule-Thomson expansion-is particularly noteworthy. Investigating the deeper physical implications of this phenomenon is crucial, and a simplified analysis is presented below.
By equating the Hawking temperature Eq.(7) with the inversion temperature Eq.(53), we obtain
$\frac{\dot{R}_{A}}{HR_{A}}=\frac{2(R_{A}^{2}+2\widetilde{\alpha})}{(n-2)R_{A}^{2}+2(n-4)\widetilde{\alpha}}\,$, (55)
which together with (5) results to the following relation
$\dot{H}-\frac{k}{a^{2}}=-\frac{2(R_{A}^{2}+2\widetilde{\alpha})}{R_{A}^{2}[(n-2)R_{A}^{2}+2(n-4)\widetilde{\alpha}]}\,$. (56)
Furthermore, by utilizing Eqs.(30) and (29), we ultimately derive the constraints on the perfect fluid in the FRW universe within the framework of Einstein-Gauss-Bonnet gravity as follows
$1+\frac{p}{\rho}=\frac{4(R_{A}^{2}+2\widetilde{\alpha})^{2}}{(n-1)(R_{A}^{2}+\widetilde{\alpha})[(n-2)R_{A}^{2}+2(n-4)\widetilde{\alpha}]}\,$. (57)
When $\alpha=0$ (or $\widetilde{\alpha}=0$ (i.e., in the absence of the Gauss-Bonnet term), we also derive the constraints on the perfect fluid in the FRW universe within the framework of Einstein gravity as follows
$1+\frac{p}{\rho}=\frac{4}{(n-1)(n-2)}\,.$ (58)
Specifically, when $n=4$, the expression (58) reduces to $\ddot{a}=0$ at the inversion point, indicating a transition between acceleration ($\ddot{a}>0$) and deceleration ($\ddot{a}<0$), or vice versa Abdusattar:2021wfv . These observations highlights the noticeable influence of spacetime dimensions on the perfect fluid and the evolution of the FRW universe.
@@ V Conclusions and Discussions
In this paper, we have studied the thermodynamics especially the equation of state and Joule-Thomson expansion of the $n$-dimensional FRW universe with a perfect fluid. From the negative surface gravity of the FRW universe on apparent horizon, $i.e.$ $\kappa<0$, we imposed a constraint on a perfect fluid in the framework of Einstein gravity and Einstein-Gauss-Bonnet gravity respectively, as detailed in Appendix A. We have derived the equations of state for the $n$-dimensional FRW universe, where the thermodynamic pressure $P$ is defined by the work density $W$, which is a natural definition directly read out from the first law of thermodynamics. We find that the $n$-dimensional FRW universe under Einstein’s gravity with perfect fluid do not have $P$-$V$ phase transitions and critical behavior. Moreover, we find that the specific capacity of the $n$-dimensional FRW universe is always positive. This result is remarkable, since it implies that the FRW universe is stable. Under Einstein-Gauss-Bonnet gravity, the appearance of $P$-$V$ phase transitions as well as (un)stability of the FRW universe with a perfect fluid are depended on the spacetime dimension $n$ and gravitational coupling parameter $\widetilde{\alpha}$.
[…]
Notably, in the present work, we introduce the $T$-$P$ isenthalpic expansion process into a high-dimensional FRW universe within the framework of modified gravity, marking the first pioneering exploration of thermodynamic properties across arbitrary dimensions. This innovation not only establishes a theoretical foundation to support future astronomical observations–potentially providing empirical anchors for our findings-but also deepens our understanding of cosmic evolution from a thermodynamic perspective. Furthermore, a key subsequent challenge, yet one of significant import, lies in determining whether cosmological observations can detect an inversion temperature, a task that bridges our theoretical framework with observational cosmology and could further validate the thermodynamic insights we have uncovered. It would be interesting to extend the present study to higher-order modified gravity theories such as Lovelock gravity, as they may reveal intriguing phenomena. Another open question arises when considering scenarios where entropy may take on different values. In our current work, our primary focus is on the case where entropy is positive, and all our discussions and analyses are grounded on this premise. However, Refs. Clunan:2004tb ; Cvetic:2001bk ; Nojiri:2002qn highlight that entropy can be negative due to the coupling of the Gauss-Bonnet term, presenting a fascinating avenue for future research to deepen our understanding of thermodynamic properties in Gauss-Bonnet gravity. It is therefore intriguing to explore the implications that would arise if we consider cases where entropy could be negative. We will pursue these investigations in the future.
@@ VI Acknowledgment
We are deeply grateful to the reviewer for their insightful and constructive comments, which have significantly improved the quality of our work. We also thank Profs.Ya-Peng Hu and Yu-Sen An for their useful discussions. Haximjan Abdusattar is supported by National Natural Science Foundation of China (NSFC) under grant No. 12465012, Kashi University high-level talent research start-up fund project under grant No. 022024002, and Tianchi Talented Young Doctors Program of Xinjiang Uygur Autonomous Region. Shi-Bei Kong is supported by grant No.12465011 from the National Natural Science Foundation of China (NSFC) and start-up fund DHBK2023002 from East China University of Technology (ECUT). Zhong-Wen Feng is supported by the Natural Science Foundation of Sichuan Province under Grant No. 24NSFSC1618.
@@ Appendix A Negativity Condition of Surface Gravity for the Arbitrary Dimensional FRW Universe
In this appendix, we derive the negativity conditions for the surface gravity (6) in both Einstein gravity and Einstein-Gauss-Bonnet gravity. By examining Eq.(6) or (7), we deduce that a negative surface gravity ($\kappa|_{R=R_{A}}<0$) implies that the second term in these equations falls into one of the following categories: less than zero, exactly zero, or greater than zero but less than one. These conditions potentially impose constraints on the pressure and density characteristics of a perfect fluid.
Using Eqs.(4), (5), and (30), we obtain the second term in Eq.(6) as
$\delta=\frac{\dot{R}_{A}}{2HR_{A}}=-\frac{R_{A}^{2}}{2}\Big{(}\dot{H}-\frac{k}{a^{2}}\Big{)}=\frac{4\pi(\rho+p)}{n-2}\frac{R_{A}^{2}}{1+{2\widetilde{\alpha}}/{R_{A}^{2}}}\,.$ (59)
We analyze the following scenarios:
1. $\delta=0$ ($\dot{R}_{A}=0$):
This leads to $T=1/(2\pi R_{A})$. Under this condition, we obtain $\rho+p=0$.
2. $\delta<0$:
This leads to $2\pi R_{A}T>1$. From Eq.(59), we obtain
$\frac{4\pi(\rho+p)}{n-2}\frac{R_{A}^{2}}{1+{2\widetilde{\alpha}}/{R_{A}^{2}}}<0\,.$ (60)
  • When $R_{A}<\sqrt{-2\widetilde{\alpha}}$, we have $\rho+p>0$.
  • When $\widetilde{\alpha}>0$, or $R_{A}>\sqrt{-2\widetilde{\alpha}}$ for $\widetilde{\alpha}<0$, we obtain $\rho+p<0$.
3. $\delta=1$:
This leads to $T=0$. From Eq.(59), we obtain the corresponding negativity condition of surface gravity in Einstein-Gauss-Bonnet gravity
$\frac{4\pi(\rho+p)}{n-2}\frac{R_{A}^{2}}{1+{2\widetilde{\alpha}}/{R_{A}^{2}}}=1\,$, (61)
which is equivalent to
$\frac{p}{\rho}=1-\frac{4}{n-1}\Big{(}1+\frac{\widetilde{\alpha}}{R_{A}^{2}+\widetilde{\alpha}}\Big{)}\,$, (62)
where we have used the Eq.(31).
4. $0<\delta<1$:
This leads to $0<2\pi R_{A}T<1$. From Eq.(59), we obtain
$0<\frac{4\pi(\rho+p)}{n-2}\frac{R_{A}^{2}}{1+{2\widetilde{\alpha}}/{R_{A}^{2}}}<1\,$, (63)
  • When $R_{A}<\sqrt{-2\widetilde{\alpha}}$, we have
    $\frac{n-2}{4\pi R_{A}^{2}}\Big{(}1+\frac{2\widetilde{\alpha}}{R_{A}^{2}}\Big{)}<\rho+p<0\,$, (64)
    which is equivalent to
    $\frac{4}{n-1}\Big{(}1+\frac{\widetilde{\alpha}}{R_{A}^{2}+\widetilde{\alpha}}\Big{)}<1+\frac{p}{\rho}<0\,.$ (65)
  • When $\widetilde{\alpha}>0$, or $R_{A}>\sqrt{-2\widetilde{\alpha}}$ for $\widetilde{\alpha}<0$, we obtain
    $0<\rho+p<\frac{n-2}{4\pi R_{A}^{2}}\Big{(}1+\frac{2\widetilde{\alpha}}{R_{A}^{2}}\Big{)}\,$, (66)
    which is equivalent to
    $0<1+\frac{p}{\rho}<\frac{4}{n-1}\Big{(}1+\frac{\widetilde{\alpha}}{R_{A}^{2}+\widetilde{\alpha}}\Big{)}\,$. (67)
In the limit $\alpha\rightarrow 0$, the condition becomes
$-1<\frac{p}{\rho}<\frac{5-n}{n-1}\,$, (68)
which is the negativity condition for the surface gravity of the $n$-dimensional FRW universe in Einstein gravity.